Grade 12Mechanics

A particle is moving with a constant angular acceleration of 4 rad /s^2 in a circular path. At time t =0, particle was at rest. Find the time at which the magnitudes of centripetal acceleration and tangential acceleration are equal.

Aoishi
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
Dear Student,
As we realize that 
Centripetal speeding up = v^2/r 
To start with condition of movement 
ω = ωo+ αt 
As beginning rakish speed is 0 
ω = αt 
v/r = 4t 
v = 4tr. 
Extraneous increasing speed = centripetal quickening 
αr = a 
4r = v^2/r 
4r = (4tr)2/r 
4 = 16t^2 
t = ½ 
t = 0.5 sec.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)